solve and name each system question

Solve and name each system of equation 12. 4x + 3y =-13 -x + y = 5 13. 7x + 2y = 6 14x + 4y 12 14. 9x-5y 1 -18x + 10y = 1 15. 2x + y + z = 9 -x-y+ z 1 3x-y +z 9
Answer #1

solve by matrix method (all)

4x+3y=-13

-x+y=5

matrix form is

4 3 -13
-1 1 5

Divide row1 by 4

1 3/4 -13/4
-1 1 5


Add (1 * row1) to row2

1 3/4 -13/4
0 7/4 7/4


Divide row2 by 7/4

1 3/4 -13/4
0 1 1


Add (-3/4 * row2) to row1

1 0 -4
0 1 1

solution is \large {\color{Red} (x,y)=(-4,1)}

.

.

7x+2y=6

14x+4y=12

matrix form is

7 2 6
14 4 12

Divide row1 by 7

1 2/7 6/7
14 4 12


Add (-14 * row1) to row2

1 2/7 6/7
0 0 0

last row is zero so system has infinitely many solution...because both line are same

2 0...........and \large y=y

so general solution is

(ע 0

.

.

.

9x-5y=1

-18x+10y=1

matrix form is

9 -5 1
-18 10 1

Divide row1 by 9

1 -5/9 1/9
-18 10 1


Add (18 * row1) to row2

1 -5/9 1/9
0 0 3


Divide row2 by 3

1 -5/9 1/9
0 0 1


Add (-1/9 * row2) to row1

1 -5/9 0
0 0 1

from last row 0x+0y=1....wich is not possible so system has no solution beacause both line is parallel

.

.

.

2x+y+z=9

-x-y+z=1

3z-y+z=9

matrix form is

2 1 1 9
-1 -1 1 1
3 -1 1 9

Divide row1 by 2

1 1/2 1/2 9/2
-1 -1 1 1
3 -1 1 9


Add (1 * row1) to row2

1 1/2 1/2 9/2
0 -1/2 3/2 11/2
3 -1 1 9


Add (-3 * row1) to row3

1 1/2 1/2 9/2
0 -1/2 3/2 11/2
0 -5/2 -1/2 -9/2


Divide row2 by -1/2

1 1/2 1/2 9/2
0 1 -3 -11
0 -5/2 -1/2 -9/2


Add (5/2 * row2) to row3

1 1/2 1/2 9/2
0 1 -3 -11
0 0 -8 -32


Divide row3 by -8

1 1/2 1/2 9/2
0 1 -3 -11
0 0 1 4


Add (3 * row3) to row2

1 1/2 1/2 9/2
0 1 0 1
0 0 1 4


Add (-1/2 * row3) to row1

1 1/2 0 5/2
0 1 0 1
0 0 1 4


Add (-1/2 * row2) to row1

1 0 0 2
0 1 0 1
0 0 1 4

solution is \large {\color{Red} (x,y,z)=(2,1,4)}